# DaRec: A Disentangled Alignment Framework for Large Language Model and Recommender System

### v1

@@ I Introduction
Recommender systems have become a hot spot recently, which play a crucial role in various applications, such as video streaming, social media, and e-commerce. Owing to the strong representation learning ability, deep neural network-based recommendation algorithms [1, 2, 3, 4] have demonstrated impressive capabilities. More recently, large language models (LLMs) have exhibited strong reasonable proficiency in many tasks, e.g., vision task [5, 6], natural language processing [7, 8], and graph [9]. Several works explore the application of LLMs in recommendation tasks, including semantic representation alignment [10, 11, 12, 13, 14, 15], representation augmentation [16, 17, 18], ranking function [19, 20, 21], etc.
@@ III-A Motivation
Theorem 1 indicates that the optimal recommendation error with the exactly aligned representations is at least $\Delta p$ larger than we can obtain from the input data if the information gap between collaborative models and LLMs is large. Furthermore, since LLMs and collaborative models have different semantic scenarios and training procedures, there is specific information for each model. Performing exact alignment with all representations will introduce the specific information of collaborative models and LLMs. This specific information may be mutual interference, leading to the downstream recommendation tasks performance decreasing. Therefore, in this paper, we first disentangle the initial representations in both the collaborative model and LLM into specific representation and shared representation. Then, we design a structure alignment strategy at both local and global levels to perform a more slack alignment. We provide the proof in section.VIII.
@@ III-D Optimization and Complexity
TABLE III: Recommendation Performance on three datasets with six metrics. The best results are denoted in bold. $\dagger$ denotes results are statistically significant where the p-value is less than 0.5.
@@ IV Theoretical analysis
We provide the proof in section IX.
@@ V-A Experimental Settings
Benchmark Datasets. The experimental results are evaluated in three widely used benchmark datasets, including Amazon Book [^1] https://cseweb.ucsd.edu/ jmcauley/datasets/amazon/links.html, Yelp [^2] https://www.yelp.com/dataset, and Steam  [^3] https://github.com/kang205/SASRec.
  • The Amazon Book dataset contains user ratings and reviews for books. It includes user IDs, item IDs, ratings, and detailed reviews. Additionally, it provides rich metadata such as book titles, authors, categories, descriptions, and cover images. This dataset is extensively used for developing and evaluating recommendation algorithms and for conducting user behavior analysis, offering a comprehensive resource for real-world data. It contains 11,000 users, 9,332 items, and 120,464 interactions.
  • The Yelp dataset is a user-business dataset that provides extensive textual and categorical information. It includes user reviews, business details, ratings, and metadata such as business categories, locations, and attributes.There are 11,091 users, 11,010 items, and 166,620 interactions in Yelp. It serves as a rich resource for studying user behavior and business interactions in real-world scenarios.
  • The Steam dataset encompasses textual feedback from users on the Steam platform, featuring contributions from 23,310 users and covering 5,237 items, resulting in a total of 316,190 interactions. It includes user reviews, ratings, and diverse metadata related to games, such as titles, genres, and release dates. This dataset is extensively utilized in academic research for the development of recommendation systems, sentiment analysis, and various other machine learning applications. It offers a valuable resource for investigating user behavior and preferences within the gaming context, facilitating a deeper understanding of user engagement and interaction dynamics in digital entertainment.
@@
Training Details. The experiments are conducted on the PyTorch deep learning platform with the 32G V100. For the baselines, we adopt their source with original settings. In our model, the learning rate is set to 1e-3 for all datasets and baselines with Adam optimizer. Following RLMRec[10], we combine the system prompt and the user/item profile to generate the prompt. Moreover, we utilize the GPT-3.5-turbo and text-embedding-ada-002 [38] to generate the representations $\textbf{E}^{\textbf{L}}$. Moreover, we set the trade-off hyper-parameter $\lambda$ as $0.1$ for all datasets and baselines. The sampling number $\hat{N}$ is set to 4096 for all experiments. The code of our DaRec is released in  [^4] https://anonymous.4open.science/r/DaRec-AB29..
@@ V-D1 Sensitivity Analysis of Cluster Number $K$
LightGCN-Amazon
LightGCN-Yelp
Figure 7: Sensitive analysis for the sampling number $\hat{N}$.
@@ V-E Visualization Analysis (RQ4)
VI Related Work
VI-A GNN-based Recommendation
Within the realm of recommender systems, collaborative filtering stands as a cornerstone technology, exerting a significant influence on the operation of these systems. Moreover, Graph Neural Networks (GNNs) have become a research spot recently [40, 41, 42, 43]. Existing methods always utilize GNNs, such as LightGCN [32], NGCF [29] and GCCF [31], to model the historical user-item interactions, thereby facilitating the capture of more complex relationships. Nonetheless, the implicit feedback data from users frequently contains considerable noise, which can compromise the performance of these Graph Neural Network (GNN)-based methods. In response to the aforementioned challenges, a self-supervised learning method, commonly referred to as contrastive learning [44, 45, 46, 47], takes precedence. Representative approaches, such as SGL [33], LightGCL [48], and NCL [49], employ the contrastive augmented data to boost the robustness of the whole recommendations and take out more promising performance.
VI-B Large Language Models
@@ VII Conclusion
VIII Proof of Theorem 1
@@
${\inf}_{h}{\mathbb{E}_{p}}[\ell_{CE}(h(\textbf{E}^{\textbf{C}},\textbf{E}^{\textbf{L}}),Y)]-{\inf}_{h^{\prime}}{\mathbb{E}_{p}}[\ell_{CE}(h^{\prime}(\textbf{E}^{\textbf{C}},\textbf{E}^{\textbf{L}}),Y)]$ (18)
@@ IX Proof of Theorem 2
where $\textbf{J}_{11}^{\textbf{D}}\in\mathbb{R}^{N\times N}$, $\textbf{J}_{12}^{\textbf{D}}\mathbb{R}^{N\times N}$, $\textbf{J}_{21}^{\textbf{D}}\in\mathbb{R}^{n\times N}$ and $\textbf{J}_{22}^{\textbf{D}}\in\mathbb{R}^{n\times n}$. $i,j\in[1,N]$ and $k,l\in[1,n]$. After that, we only have to proof $\textbf{J}_{12}$ is an all-zero matrix while the determinant of $\textbf{J}_{11}^{\textbf{D}}$ is non-zero to show that the matrix consisting of all the partial derivatives of $f_{sh}^{\textbf{D}}(\cdot)$ w.r.t. $\textbf{E}_{sh}^{*}$ is full rank while any partial derivatives of $f_{sh}^{\textbf{D}}(\cdot)$ w.r.t. $\textbf{E}_{sp}^{\textbf{D}*}$ is zero.
With any fixed $\bar{\textbf{E}}_{sh}^{*}$ and $\bar{\textbf{E}}_{sp}^{\textbf{D'}*}$, for all $\textbf{E}_{sh}^{\textbf{D}*}$, we have:
@@
$=0.$
Therefore, based on the above proof, we could obtain $H(\widehat{\textbf{E}}^{\textbf{D}})-H(\widetilde{\textbf{E}}^{\textbf{D}})=H(\widehat{\textbf{E}}^{\textbf{D}}\mid\textbf{D'})-H(\widetilde{\textbf{E}}^{\textbf{D}}\mid\textbf{D'})$’ ).
We could divide Theorem 12 into two components. We proof the first as follows. We use the complement information of the representation extracted by our designed method and previous method as $I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})$’ ) and $I(\widetilde{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})$’ ). Since we split the representations into two components and we perform the structure alignment in shared part, we have $I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})\geq I(\widetilde{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})$’ ). Thus, we have:
$I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R})=I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R},\textbf{D'})+I(\widehat{\textbf{E}},\textbf{R}|\textbf{D'}).$’ ) . (28)
@@
Moreover, $I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})\geq I(\widetilde{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})$’ ), we have $I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R})\geq I(\widetilde{\textbf{E}}^{\textbf{D}},\textbf{R})$.

### v2

@@ I Introduction
Recommender systems have become a hot spot recently, which play a crucial role in various applications, such as video streaming, social media, and e-commerce. Owing to the strong representation learning ability, deep neural network-based recommendation algorithms [1, 2, 3, 4, 5, 6, 7] have demonstrated impressive capabilities. More recently, large language models (LLMs) have exhibited strong reasonable proficiency in many tasks, e.g., vision task [8, 9], natural language processing [10, 11], and graph [12]. Several works explore the application of LLMs in recommendation tasks, including semantic representation alignment [13, 14, 15, 16, 17, 18], representation augmentation [19, 20, 21], ranking function [22, 23, 24], etc.
@@ III-A Motivation
Theorem 1 indicates that the optimal recommendation error with the exactly aligned representations is at least $\Delta p$ larger than we can obtain from the input data if the information gap between collaborative models and LLMs is large. Furthermore, since LLMs and collaborative models have different semantic scenarios and training procedures, there is specific information for each model. Performing exact alignment with all representations will introduce the specific information of collaborative models and LLMs. This specific information may be mutual interference, leading to the downstream recommendation tasks performance decreasing. Therefore, in this paper, we first disentangle the initial representations in both the collaborative model and LLM into specific representation and shared representation. Then, we design a structure alignment strategy at both local and global levels to perform a more slack alignment. We provide the proof in section.IX.
@@ III-D Optimization and Complexity
TABLE III: Recommendation Performance on three datasets with six metrics. The best results are denoted in bold. $\dagger$ denotes results are statistically significant where the p-value is less than 0.05.
@@ IV Theoretical analysis
We provide the proof in section X.
@@ V-A Experimental Settings
Benchmark Datasets. The experimental results are evaluated in three widely used benchmark datasets, including Amazon Book, Yelp, and Steam.
@@
Training Details. The experiments are conducted on the PyTorch deep learning platform with the 32G V100. For the baselines, we adopt their source with original settings. In our model, the learning rate is set to 1e-3 for all datasets and baselines with Adam optimizer. Following RLMRec[13], we combine the system prompt and the user/item profile to generate the prompt. Moreover, we utilize the GPT-3.5-turbo and text-embedding-ada-002 [41] to generate the representations $\textbf{E}^{\textbf{L}}$. Moreover, we set the trade-off hyper-parameter $\lambda$ as $0.1$ for all datasets and baselines. The sampling number $\hat{N}$ is set to 4096 for all experiments.
@@ V-D3 Sensitivity Analysis of sampling size $\hat{N}$
LightGCN-Amazon
LightGCN-Yelp
Figure 7: Sensitive analysis for the sampling number $\hat{N}$.
@@ V-E Visualization Analysis (RQ4)
VI Case Study
In this section, we conduct a case study to demonstrate the effectiveness of our DaRec framework. We explore how LLMs enhance the semantic features of collaborative models through our designed alignment framework. Specifically, we leverage the model’s ability to capture global user dependencies. We focus on users who are separated by multiple hops ($>$ 5 hops) in the network. To evaluate the model’s ability to capture these global relationships, we calculate the similarity of user representations. For this purpose, we adopt SimGCL [37], RLMRec-Con [13], and our DaRec as baselines, all employing the same backbone. The dataset used for this study is Yelp. The relationships are evaluated using two metrics: relevance score and the ranking of long-distance neighbors based on this score. The relevance score is determined using the cosine similarity function. The case study is presented in Fig. 8. In this scenario, we focus on user $u_{2734}$ and user $u_{3648}$. From the results, we observe that with our designed alignment framework DaRec, the semantic information are better aligned between $u_{2734}$ and $u_{3648}$, i.e., ”snacks” and ”diverse textures”. The relevance score and the ranking are increasing. This demonstrates that the learned representations from our DaRec capture global collaborative relationships beyond other recommendation methods.
VII Related Work
VII-A GNN-based Recommendation
Within the realm of recommender systems, collaborative filtering stands as a cornerstone technology, exerting a significant influence on the operation of these systems. Existing methods always utilize Graph Neural Networks (GNNs), such as LightGCN [35], NGCF [32] and GCCF [34], to model the historical user-item interactions, thereby facilitating the capture of more complex relationships. Nonetheless, the implicit feedback data from users frequently contains considerable noise, which can compromise the performance of these Graph Neural Network (GNN)-based methods [43, 44, 45, 46, 47, 48, 49]. In response to the aforementioned challenges, a self-supervised learning method, commonly referred to as contrastive learning, takes precedence. Representative approaches, such as SGL [36], LightGCL [50], and NCL [51], employ the contrastive augmented data to boost the robustness of the whole recommendations and take out more promising performance.
Figure 8: Case study to demonstrate the ability on capturing global user dependencies.
VII-B Large Language Models
@@ VIII Conclusion
Acknowledgment
This work was supported by the National Key R&D Program of China 2020AAA0107100, the Natural Science Foundation of China (project no. 62325604, 62276271, 62476281).
IX Proof of Theorem 1
@@
${\inf}_{h}{\mathbb{E}_{p}}[\ell_{CE}(h(\textbf{E}^{\textbf{C}},\textbf{E}^{\textbf{L}}),Y)]-{\inf}_{h^{\prime}}{\mathbb{E}_{p}}[\ell_{CE}(h^{\prime}(\textbf{E}^{\textbf{C}},\textbf{E}^{\textbf{L}}),Y)]$
@@ X Proof of Theorem 2
where $\textbf{J}_{11}^{\textbf{D}}\in\mathbb{R}^{N\times N}$, $\textbf{J}_{12}^{\textbf{D}}\mathbb{R}^{N\times N}$, $\textbf{J}_{21}^{\textbf{D}}\in\mathbb{R}^{n\times N}$ and $\textbf{J}_{22}^{\textbf{D}}\in\mathbb{R}^{n\times n}$. $i,j\in[1,N]$ and $k,l\in[1,n]$. After that, we only have to proof $\textbf{J}_{12}$ is an all-zero matrix while the determinant of $\textbf{J}_{11}^{\textbf{D}}$ is non-zero to show that the matrix consisting of all the partial derivatives of $f_{sh}^{\textbf{D}}(\cdot)$ w.r.t. $\textbf{E}_{sh}^{*}$ is full rank while any partial derivatives of $f_{sh}^{\textbf{D}}(\cdot)$ w.r.t. $\textbf{E}_{sp}^{\textbf{D}*}$ is zero. With any fixed $\bar{\textbf{E}}_{sh}^{*}$ and $\bar{\textbf{E}}_{sp}^{\textbf{D'}*}$, for all $\textbf{E}_{sh}^{\textbf{D}*}$, we have:
@@
$=0.\vspace{-5pt}$
Therefore, based on the above proof, we could obtain $H(\widehat{\textbf{E}}^{\textbf{D}})-H(\widetilde{\textbf{E}}^{\textbf{D}})=H(\widehat{\textbf{E}}^{\textbf{D}}\mid\textbf{D'})-H(\widetilde{\textbf{E}}^{\textbf{D}}\mid\textbf{D'})$’ ). We could divide Theorem 12 into two components. We proof the first as follows. We use the complement information of the representation extracted by our designed method and previous method as $I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})$’ ) and $I(\widetilde{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})$’ ). Since we split the representations into two components and we perform the structure alignment in shared part, we have $I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})\geq I(\widetilde{\textbf{E}}^{\textbf{D}},\textbf{R}|\textbf{D'})$’ ). Thus, we have $I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R})=I(\widehat{\textbf{E}}^{\textbf{D}},\textbf{R},\textbf{D'})+I(\widehat{\textbf{E}},\textbf{R}|\textbf{D'})$’ ).